In the laboratory you dilute 3.54 mL of a concentrated 12.0 M nitric acid solution to a total volume of 50.0 mL. What is the concentration of the dilute solution?

__________ M.
1
Expert's answer
2019-03-18T02:19:28-0400

Solution.

Firstly, we must find the amount of substance in the initial solution according to this formula:

n = Cm * V

n = (3.54 * 12)/1000 = 0.0425 mole

Secondly, to find the concentration of a new solution, use the following formula:

Cm = n/V(new)

Cm = 0.0425/0.05 = 0.85 M

Answer:

Cm = 0.0425/0.05 = 0.85 M